OpenStudy (anonymous):

Hi, so my teacher told me about this assignment I have to do. I gave it to him, he gave me a 50 percent and told me what to do to help me, so please help me. I will give you fan and medal :) Thanks

OpenStudy (anonymous):

Assignment Instructions Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. Stock A B C Price: f(x) = 25(1.08)x X F(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Step 1: Write equations for each Stock (fill in the blanks) Stock A B C D Values for D Starting price 25 22 30 Second lowest starting price: 23 23 Rate of growth 8% each day 2% each day 4% each day Second highest rate of growth: 6% each day 6% Equation f(x) = 25(1.08)x F(x) = 22(1.02)x F(x)= 30(1.4)x F(x) = 23(1.06)x F(x) = 23(1.06)x   Graph the price function for Stock D in Geogebra. Past your graph below: What is the y-intercept? (0,23) What does the y-intercept represent? The starting value for stock D Is this a growth function? Why? Yes, because it is constantly increasing. What is the rate of growth? 6% Explain the key features of the graph with complete sentences. A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). Hint: Does this shift the function up or down? Past your graph from Geogebra here: Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. https://www.desmos.com/calculator/0zzwz6u4gk -→ the graph for all of them Function Transformation from f(x) f(x + 2) The profit decreases. The profit is decreasing over time. f(–x) The profit decreases as x goes to the right side. The profit is decreasing over time. f(x) + 3 The profit decreases .The profit is decreasing over time.

OpenStudy (anonymous):

(It is better in the document form)

OpenStudy (anonymous):

My teacher told me that I did a great job on my equation for the 1st part. The graph i attached was for number 4 and I need a graph for number 2, he told that it should only show your equation. For number 3 it should have 2 functions , my function and then f(X) _ g(X). Thank you for anybody who helps me @hartnn @TheRaggedyDoctor

OpenStudy (anonymous):

@TheSmartOne please help